Please help me with this problem, Thanks
Use the Completing the square or shortcut Method to find the Standard Form of the following quadratic function. Standard form: f(x)=a(x-h)^2+k
1. f(x)=x^2-2x-5=_______

Answer by uma(370)

f(x) = x^2 - 2x - 5
==> f(x) = x^2 - 2x + 1 - 1 - 5 [adding and subtracting 4 to complete the square]
==> f(x) = x^2 - 2x + 1 - 6
= (x^2 - 2x + 1) - 6
= (X-1)^2 - 6
This is of the standard form f(x) = a(x-h)^2 + k where,
a = 1; h = 1; k = -6


Good Luck!!!